# FlexiClip: Locality-Preserving Free-Form Character Animation

### v1

@@ 4.1 Experimental Setup
FlexiClip enables high-resolution, complex animations by leveraging SVGs, where paths defined by control points are animated via mesh deformation and rendered into bitmaps using DiffVG (Li et al., 2020) for video SDS loss. For bitmap clipart, pixels within each triangle are warped. We used 30 clipart images from AniClipart (Wu et al., 2024) and additional ones from Freepik [^1] https://www.freepik.com/ across various categories (humans, animals, and objects), resized to 256×256 pixels. First, we learn $\mathbf{J}_{0}$ with Eq.4 for 10K iters. After that, motion trajectories with 8–11 control points were optimized upto 700 steps using Adam (learning rate: 0.5). We applied ModelScope T2V (Wang et al., 2023) with a guidance parameter of 50 for SDS loss. For spatial posing, cubic Bézier control points, we use a 4-layer MLP with LeakyReLU activation, with the final layer being linear. Temporal Jacobians were predicted with a 3-layer MLP, while two attention networks with 32-dimensional keys/values and two heads modeled motion and deformation effectively. Standard 24-frame animations were rendered on an NVIDIA V100 in  40 minutes using 26 GB.
@@ 4.2 Metrics
Deformation Smoothness(DS): This metric evaluates how smoothly the control points deform along Bézier paths. It computes the average difference in displacement vectors of consecutive frames:
@@
Animation Energy(AE): To evaluate the energy distributed across control points, we calculate the mean squared displacement across all frames:
@@ 4.3 Comparison to State-of-the-Art Methods
Figure 3: FlexiClip vs AniClipart: Four consecutive frames are shown for comparison. AniClipart distorts the objects (e.g., boy/girl jumping, woman dancing), lacks proper text conditioning (e.g., parrot), and shows poor temporal consistency (e.g., boy jumping). In contrast, FlexiClip preserves the visual identity, maintaining consistent shapes and producing high-quality, smooth, and text-aligned animations. Detailed animations can be seen at: https://flexiclip.github.io/
@@
Qualitatively, we compare FlexiClip with AniClipart based on the results observed in the tables and visual examples. One of the key differences between the two models lies in their ability to preserve visual identity and create more coherent text-video animations. While AniClipart demonstrates impressive performance in visual identity preservation and text-video alignment, it encounters challenges when it comes to handling complex deformations. For instance, in the case of hand deformation (e.g., during dance movements), AniClipart exhibits noticeable distortions, which negatively impact the naturalness of the animation. FlexiClip, on the other hand, maintains better consistency in object shape and deformation, showing smoother transitions without compromising visual identity. In particular, when handling leg folding during jumping and dynamic hand movements, FlexiClip ensures these actions are portrayed realistically (whereas AniClipart shows the distorted hand girl/boy jumping), which is a critical requirement for creating believable animations.
@@
Additionally, FlexiClip excels at producing smoother and more realistic deformations, as evidenced by the parrot animation, where the wing flapping looks much more natural and consistent. The wings of the parrot exhibit a smooth, continuous motion, enhancing the visual realism compared to AniClipart’s deformation, which is flapping its tail rather than wings showing lack of text alignment. Looking at the quantitative results (Tab.1 and 2), FlexiClip outperforms AniClipart in both the bitmap and animation/vector metrics. Specifically, FlexiClip achieves a higher CLIP score (0.9563 vs. 0.9401) and X-CLIP score (0.2102 vs. 0.2075), signifying better visual identity preservation and text-video alignment. Furthermore, FlexiClip’s animation quality metrics, including MV, TC, GD, DS, and AE, also surpass AniClipart’s results, with FlexiClip yielding more motion variation, lower distortions, and higher animation efficacy. Notably FlexiClip GD has increased but DS decreased showing smoother animations. FlexiClip versus I2V Models: See App.B User Study: To assess the improvements made by FlexiClip, we conducted a subjective user study with 55 static clipart images animated by six methods: FlexiClip (Ours), AniClipart, LTXVideo, PyramidFlow, AnimateLCM-I2V, and DynamiCrafter. Participants were asked to rate the animations on visual identity preservation, text-video alignment, and smoothness using a six-point scale from 0 (strongly disagree) to 1.0 (strongly agree). The study was conducted online with 30 participants, and ratings were averaged across the 55 clipart images. The results in Tab.3 show that FlexiClip significantly outperforms the other methods.
@@ 4.4 Ablation Study
We conducted an ablation study to validate the significance of each critical component in FlexiClip. The quantitative results of the study are detailed in Table 4, while Figure 7 and 7 showcases a qualitative comparison of the different variants. Below, we describe the importance of temporal Jacobian obtained from pfODE and flow matching loss, which constitute the core of our system. Notably, the geometric deviation (GD) in FlexiClip is higher compared to AniClipart due to the absence of ARAP deformation, which inherently minimizes shape distortion.
Temporal Jacobian. To evaluate the effectiveness of incorporating temporal Jacobian, we removed this component and instead used a baseline keypoint transformation without considering the continuous dynamics provided by the pfODE framework. Without temporal Jacobian, the animation exhibited noticeable artifacts (Fig.7), including a lack of smoothness in movements and unnatural keypoint transitions, as seen in the parrot, ghost, and cloud examples in Fig.7. The quantitative results corroborate these findings: the ”w/o Temporal Jacobian” variant exhibited a lower motion variance (MV = 23.00) and higher temporal inconsistency (TC = 8.80). Additionally, the geometric distortion (GD) score for this variant decreased to 51.50 (vs Default) showing rigid transformation as seen from ghost, parrot example it failed to show movement of hands and wings flapping respectively. In contrast, the default setting, with temporal Jacobians, yielded superior results across all metrics, highlighting its necessity for realistic and accurate animations.
@@
Flow Matching Loss. The flow matching loss plays a crucial role in aligning the generated motion trajectories with the temporal coherence and dynamics inferred from text descriptions. To assess its impact, we omitted the flow matching loss from the total loss. As shown in the man and woman dance examples in Fig.7, this substitution led to erratic and exaggerated movements, deviating significantly from the intended motion semantics. Quantitatively, the ”w/o Flow Match. Loss” variant showed increased geometric distortion (GD = 53.00) and reduced average energy (AE = 95.00), indicative of less expressive and abrupt motion dynamics. Additionally, temporal consistency (TC) degraded slightly (TC = 8.40), reinforcing the importance of this loss in maintaining stable frame-to-frame transitions. The absence of flow matching loss caused inconsistencies visible in the chaotic limb movements and unstable/abrupt posture transitions in the dancing examples. The default setting, which incorporates flow matching loss, demonstrated a balanced trade-off between dynamism and coherence, achieving the highest average energy (AE = 113.44) and a competitive motion variance (MV = 25.33) with smooth animation.
@@ 4.5 More Results
Rotation: Demonstrated rotational movements. The flower swaying its petals in the breeze showcases FlexiClip’s ability to handle smooth rotational dynamics. Without any rotational mechanism explicitly, the system captures the gentle oscillation of petals, creating a lifelike representation of natural motion.
Multiple Text Conditions: Demonstrated complex motions via multiple text conditioning. The woman in a green dress with black polka dots demonstrates FlexiClip’s capability to animate based on complex textual conditions. Here, she dances and folds her hands in rhythm with the description, reflecting the system’s ability to integrate fine-tuned motion semantics with precise keypoint dynamics.
Multiple Objects: Demonstrated coordinated interactions. The couple dancing exemplifies FlexiClip’s support for animating multiple objects simultaneously. The synchronized movements of the man and woman highlight the system’s proficiency in managing interactions between objects while maintaining spatial and temporal coherence.
Layered Animations: FlexiClip supports layered animations (see in Fig.1 Breakdancer animation) for depth and complexity. This allows for the creation of dynamic animations enhancing its applicability for interactive media.

### v2

@@ 4.1 Experimental Setup
FlexiClip enables high-resolution, complex animations by leveraging SVGs, where paths defined by control points are animated via mesh deformation and rendered into bitmaps using DiffVG (Li et al., 2020) for video SDS loss. For bitmap clipart, pixels within each triangle are warped. We used 30 clipart images from AniClipart (Wu et al., 2024) and additional ones from Freepik [^1] https://www.freepik.com/ across various categories (humans, animals, and objects), resized to 256×256 pixels. First, we learn $\mathbf{J}_{0}$ with Eq.4 for 10K iters. After that, motion trajectories with 8–11 control points were optimized upto 700 steps using Adam (learning rate: 0.5). We applied ModelScope T2V (Wang et al., 2023) with a guidance parameter of 50 for SDS loss. For spatial posing, cubic Bézier control points, we use a 4-layer MLP with LeakyReLU activation, with the final layer being linear. Temporal Jacobians from pfODE were predicted with a 3-layer MLP, while two attention networks with 32-dimensional keys/values and two heads modeled motion and deformation effectively. Standard 24-frame animations were rendered on an NVIDIA V100 in  40 minutes using 26 GB.
@@ 4.2 Metrics
Deformation Smoothness(DS):
This metric evaluates how smoothly the control points deform along Bézier paths. It computes the average difference in displacement vectors of consecutive frames:
@@
Animation Energy(AE):
To evaluate the energy distributed across control points, we calculate the mean squared displacement across all frames:
@@ 4.3 Comparison to State-of-the-Art Methods
Figure 3: FlexiClip vs AniClipart: Four consecutive frames are shown for comparison. AniClipart distorts the objects (e.g., boy/girl jumping, woman dancing), lacks proper text conditioning (e.g., parrot), and shows poor temporal consistency (e.g., boy jumping). In contrast, FlexiClip preserves the visual identity, maintaining consistent shapes and producing high-quality, smooth, and text-aligned animations. Detailed animations can be seen at: https://creative-gen.github.io/flexiclip.github.io/
@@
We compare FlexiClip with AniClipart based on the results observed in the tables (Tab.1, 2 & 3) and visual examples (Fig.3). One of the key differences between the two models lies in their ability to preserve visual identity and create more coherent text-video animations. While AniClipart demonstrates impressive performance in visual identity preservation and text-video alignment, it encounters challenges when it comes to handling complex deformations. For instance (Fig.3), in the case of hand deformation (e.g., during dance movements), AniClipart exhibits noticeable distortions, which negatively impact the naturalness of the animation. FlexiClip, on the other hand, maintains better consistency in object shape and deformation, showing smoother transitions without compromising visual identity. In particular, when handling leg folding during jumping and dynamic hand movements, FlexiClip ensures these actions are portrayed realistically (whereas AniClipart shows the distorted hand for girl/boy jumping), which is a critical requirement for creating natural animations.
@@
Additionally, FlexiClip excels at producing smoother and more realistic deformations, as evidenced by the parrot animation, where the wing flapping looks much more natural and consistent. The wings of the parrot exhibit a smooth, continuous motion, enhancing the visual realism compared to AniClipart’s deformation, which is flapping its tail rather than wings showing lack of text alignment. Looking at the quantitative results (Tab.1 and 2), FlexiClip outperforms AniClipart in both the bitmap and animation metrics. Specifically, FlexiClip achieves a higher CLIP score (0.9563 vs. 0.9401) and X-CLIP score (0.2102 vs. 0.2075), signifying better visual identity preservation and text-video alignment. Furthermore, FlexiClip’s animation quality metrics, including MV, TC, GD, DS, and AE, also surpass AniClipart’s results, with FlexiClip yielding more motion variation, lower distortions, and higher animation efficacy. Notably FlexiClip GD has increased but DS decreased showing smoother animations. FlexiClip versus I2V Models: See App.B.
User Study
To assess the improvements made by FlexiClip, we conducted a subjective user study with 55 static clipart images animated by six methods: FlexiClip (Ours), AniClipart, LTXVideo, PyramidFlow, AnimateLCM-I2V, and DynamiCrafter. Participants were asked to rate the animations on visual identity preservation, text-video alignment, and smoothness using a six-point scale from 0 (strongly disagree) to 1.0 (strongly agree). The study was conducted online with 30 participants, and ratings were averaged across the 55 clipart images. The results in Tab.3 show that FlexiClip significantly outperforms the other methods.
@@
4.4 Effect of $\lambda$ on Motion Quality and Convergence
In our experiments, we found that the $\lambda$ parameter plays a critical role in balancing the alignment of generated motion with the text prompt and ensuring smooth, natural motion trajectories. Specifically, tuning $\lambda$ influences both the quality of the animation and the convergence behavior during training. We set $\lambda=15$ for all reported results to maintain consistency across examples.
When $\lambda$ is set too low, such as around 1, the model struggles to generate coherent motion aligned with the textual description, often requiring up to 1000 gradient descent steps to produce acceptable outputs. Increasing $\lambda$ incrementally to values up to 5 does not result in a significant reduction in the required number of training steps, indicating limited benefit in that range. However, a more noticeable impact emerges when $\lambda$ is raised to values between 5 and 10, where we observe variability in convergence, some animations achieve satisfactory quality in as few as 600 steps, while others still require up to 900 steps.
At $\lambda=15$, the motion generation becomes consistently more reliable, with most animations aligning well with the text prompt after approximately 700 training steps. This value represents a practical trade-off between convergence speed and motion quality. Pushing $\lambda$ beyond 15 leads to noticeably sharper and faster motions, but this comes at the expense of temporal smoothness, resulting in less natural-looking animations. Thus, $\lambda=15$ serves as an effective default for balancing alignment, smoothness, and training efficiency.
@@ 4.5 Ablation Study
We conducted an ablation study to validate the significance of each critical component in FlexiClip. The quantitative results of the study are detailed in Table 4, while Figure 7 and 7 showcases a qualitative comparison of the different variants. We performed ablation to describe importance of temporal Jacobian obtained from pfODE and the flow matching loss, which constitute the core of our system. Notably, the geometric deviation (GD) in FlexiClip is higher compared to AniClipart due to the absence of ARAP deformation, which inherently minimizes shape distortion observed from reduced deformation smoothness(DS).
Temporal Jacobian
To evaluate the effectiveness of incorporating temporal Jacobian, we removed this component and instead used a baseline keypoint transformation without considering the continuous dynamics provided by the pfODE framework. Without temporal Jacobian, the animation exhibited noticeable artifacts (Fig.7), including a lack of smoothness in movements and unnatural keypoint transitions, as seen in the parrot, ghost, and cloud examples in (Fig.7). The quantitative results corroborate these findings: the ”w/o Temporal Jacobian” variant exhibited a lower motion variance (MV = 23.00) and higher temporal inconsistency (TC = 8.80). Additionally, the geometric distortion (GD) score for this variant decreased to 51.50 (vs Default) showing rigid transformation as seen from ghost, parrot example it failed to show movement of hands and wings flapping respectively. In contrast, the default setting, with temporal Jacobians, yielded superior results across all metrics, highlighting its necessity for realistic and accurate animations.
@@
Flow Matching Loss
The flow matching loss plays a crucial role in aligning the generated motion trajectories with the temporal coherence and dynamics inferred from text descriptions. To assess its impact, we omitted the flow matching loss from the total loss. As shown in the man and woman dance examples in Fig.7, this substitution led to erratic and exaggerated movements, deviating significantly from the intended motion semantics. Quantitatively, ”w/o Flow Match. Loss” variant showed increased geometric distortion (GD = 53.00) and reduced average energy (AE = 95.00), indicative of less expressive and abrupt motion dynamics. Additionally, temporal consistency (TC) degraded slightly (TC = 8.40), reinforcing the importance of this loss in maintaining stable frame-to-frame transitions. The absence of flow matching loss caused inconsistencies visible in the chaotic limb movements and unstable/abrupt posture transitions in the dancing examples. The default setting, which incorporates flow matching loss, demonstrated a balanced trade-off between dynamism and coherence, achieving the highest average energy (AE = 113.44) and a competitive motion variance (MV = 25.33) with smooth animation.
@@ 4.6 More Results
Rotation
Demonstrated rotational movements. The flower swaying its petals in the breeze showcases FlexiClip’s ability to handle smooth rotational dynamics. Without any rotational mechanism explicitly, the system captures the gentle oscillation of petals, creating a lifelike representation of natural motion.
Multiple Text Conditions
Demonstrated complex motions via multiple text conditioning. The woman in a green dress with black polka dots demonstrates FlexiClip’s capability to animate based on complex textual conditions. Here, she dances and folds her hands in rhythm with the description, reflecting the system’s ability to integrate fine-tuned motion semantics with precise keypoint dynamics.
Multiple Objects
Demonstrated coordinated interactions. The couple dancing exemplifies FlexiClip’s support for animating multiple objects simultaneously. The synchronized movements of the man and woman highlight the system’s proficiency in managing interactions between objects while maintaining spatial and temporal coherence.
Layered Animations
FlexiClip supports layered animations (see in Fig.1 Breakdancer animation) for depth and complexity. This allows for the creation of dynamic animations enhancing its applicability for interactive media.
@@ 5 Conclusion
Impact Statement
This paper introduces FlexiClip, a novel framework for animating static clipart images with enhanced temporal coherence and geometric integrity. By integrating temporal Jacobians, continuous-time modeling via probability flow ODEs (pfODEs), and a flow matching loss inspired by GFlowNet principles, FlexiClip addresses longstanding challenges in clipart animation, such as abrupt motions and geometric distortions.
Potential Positive Impacts
  • Advancement in Digital Animation: FlexiClip’s methodology can significantly benefit industries involved in digital animation, including education, entertainment, and digital marketing, by enabling the creation of smooth and natural animations from static images.
  • Accessibility for Content Creators: By simplifying the animation process for clipart images, FlexiClip can empower individual creators and small businesses to produce high-quality animations without extensive resources.
  • Educational Tools Enhancement: The ability to animate educational clipart can lead to more engaging learning materials, potentially improving educational outcomes.
Potential Negative Impacts and Mitigations
  • Misinformation Risks: The ease of animating static images could be misused to create deceptive content. To mitigate this, it’s essential to develop and integrate detection tools that can identify AI-generated animations.
  • Intellectual Property Concerns: Animating copyrighted clipart without proper authorization could infringe on intellectual property rights. Users should be educated on copyright laws, and systems should include checks to prevent unauthorized use.
  • Bias in Animation Outputs: If the training data for FlexiClip contains biased representations, the animations produced might perpetuate these biases. It’s crucial to use diverse and representative datasets and implement bias detection mechanisms.
Ethical Considerations
While FlexiClip primarily aims to advance the field of machine learning and animation, it is important to remain vigilant about its applications. Developers and users should adhere to ethical guidelines, ensuring that the technology is used responsibly and does not contribute to societal harm.